Culver Corporation purchased a piece of equipment for $69,900. It estimated a 6-year life and $3,300 salvage value. At the end of year 3 (before the depreciation adjustment), it estimated the new total life to be 8 years and the new salvage value to be $4,400. Compute the revised depreciation. Company uses straight-line depreciation method. (Round answer to 0 decimal places, e.g. 125.)
Answer #1
Annual depreciation for first 2 years=(69900-3300)/6*2= 22200
Book value after year 2=69900-22200= $47700
Revised depreciation=(47700-4400)/6= $7217